1. 3x² + 2x – 1=0
Ans.

2. x³-27
Ans.

State the possible rational zeros for each polynomial. Then find all rational zeros.
3. P(x) x³ +4x² + 5x + 2
Possible zeros:
Rational zeros:

4. 3x³ + 11x² + 5x – 3
Possible zeros:
Rational zeros:
1
Expert's answer
2020-11-23T05:23:37-0500

1.

"3x\u00b2 + 2x \u2013 1=0"

Here, we can apply quadratic formula

"x = \frac{−b± \sqrt{b^2−4ac} }{2a}"

Where a = 3, b = 2, c = -1

Plugging these into the equation, we have;

"x = \frac{−2± \sqrt{2^2−(4×3×-1)} }{2×-1}"


"x = \frac{−2±√16}{6}"


"x= \frac{1}{3}" or "x = -1"


2.

"x^3 - 27 =0" -27 crosses the other side

"\implies x^3 = 27" take the cube root of both sides

"\implies \sqrt[3]{x^3} = \sqrt[3]{27}"

"x = 3"


3.

"\mathrm{For\:a\:polynomial\:equation\:with\:integer\:coefficients:\quad }a_nx^n+a_{n-1}x^{n-1}+\ldots +a_0"

"\mathrm{If\:}a_0\mathrm{\:and\:}a_n\mathrm{\:are\:integers,\:then\:if\:there\:is\:a\:rational\:solution}"


"it\:could\:be\:found\:by\:checking\:all\:the\:numbers\:produced"

"\:for\:\pm \frac{dividers\:of\:a_0}{dividers\:of\:a_n}"

Therefore, given the polynomial x³ +4x² + 5x + 2, we have;

"a_0=2, and\:\quad a_n=1"

"\mathrm{The\:dividers\:of\:}a_0:\quad 1,\:2, and\:\quad \mathrm{The\:dividers\:of\:}a_n:\quad 1"

So, "\mathrm{The\:possible\:zeros\:numbers\:are\:\:}\quad \pm \frac{1,\:2}{1}" or "\quad \pm \frac{1}{1}" , "\quad \pm \frac{2}{1}"


And the rational zeros are obtained when these are replaced in the polynomial "x^3+4x^2+5x+2" , we have "x = -1, and , x=-2"


4.

"3x\u00b3 + 11x\u00b2 + 5x \u2013 3"

The above procedure applies to this polynomial where;

"a_0=3, and\:\quad a_n=3"

And "\mathrm{the\:dividers\:of\:}a_0:\quad 1,\:3,\:\quad \mathrm{The\:dividers\:of\:}a_n:\quad 1,\:3"

Hence the possible zeros are "\pm \frac{1,\:3}{1,\:3}"

And the rational zeros are, when plugged back into the polynomial

"3x\u00b3 + 11x\u00b2 + 5x \u2013 3" , we have; "x=-1,\:x=-3,\:x=\frac{1}{3}"